# Everything everywhere all at once, a probability-based enhanced sampling approach to rare events

### v1

@@ Abstract
The problem of studying rare events is central to many areas of computer simulations. In a recent paper [Kang, P., et al., Nat. Comput. Sci. 4, 451–460, 2024], Kang, Trizio, and Parrinello (2024) we have shown that a powerful way of solving this problem passes through the computation of the committor function, and we have demonstrated how the committor can be iteratively computed in a variational way and the transition state ensemble efficiently sampled. Here, we greatly ameliorate this procedure by combining it with a metadynamics-like enhanced sampling approach in which a logarithmic function of the committor is used as a collective variable. This integrated procedure leads to an accurate and balanced sampling of the free energy surface in which transition states and metastable basins are studied with the same thoroughness. We also show that our approach can be used in cases in which competing reactive paths are possible and intermediate metastable are encountered. In addition, we demonstrate how physical insights can be obtained from the optimized committor model and the sampled data, thus providing a full characterization of the rare event under study. We ascribe the success of this approach to the use of a probability-based description of rare events.
@@ I Introduction
In the last few decades, our group has contributed to this collective effort and very recently has proposed a new practical method to solve the rare event problem, which is based on computing the committor function $q(\mathbf{x})$ and using it as a sampling tool, which we summarize in Sec. III.1Kang, Trizio, and Parrinello (2024) We recall here that the committor is a function of the atomic coordinates $\mathbf{x}$ which, given two metastable states $A$ and $B$, is defined as the probability that a trajectory started at $\mathbf{x}$ ends in $B$ without having first passed by $A$Weinan and Vanden-Eijnden (2010) The committor function is arguably the most precisely defined way of describing rare events and is a quantity that remains well-defined even if the transition from $A$ to $B$ follows different competing pathways. As the committor has so many attractive features, much work has been done to determine it. Bolhuis, Dellago, and Chandler (2000); Ma and Dinner (2005); Berezhkovskii and Szabo (2005); Peters and Trout (2006); E. and Vanden-Eijnden (2006); Rotskoff, Mitchell, and Vanden-Eijnden (2022); Chen et al. (2023); Chen, Roux, and Chipot (2023); Li, Lin, and Ren (2019); Jung et al. (2023); Mitchell and Rotskoff (2024); He, Chipot, and Roux (2022); Khoo, Lu, and Ying (2019)
@@
where the average is over the Boltzmann distribution associated with the interatomic potential $U(\mathbf{x})$ at inverse temperature $\beta$, and $\nabla_{\mathbf{u}}$ denotes the gradient with respect to the mass-weighted coordinates.
In Ref. 1, we computed the variational minimum of $K[q(\mathbf{x})]$ using a self-consistent iterative procedure that to be started, only requires that the initial and final states are known. Key to the success of this approach has been the use of a bias potential that is a functional of the committor
@@
which is able to focus sampling on the otherwise elusive transition state region (see Fig. 1). This approach has proven to be highly successful in finding and extensively sampling the transition state ensemble (TSE). However, additional simulations were needed to evaluate the system free energy since $V_{K}(\mathbf{x})$ disproportionately favors sampling of the transition state ensemble region. This additional calculation can be performed, for instance, using the data collected during the iterative process to generate an efficient collective variable Bonati, Piccini, and Parrinello (2021); Ray, Trizio, and Parrinello (2023) (CV) and then performing a metadynamics-like calculation using this CV. Laio and Parrinello (2002); Invernizzi and Parrinello (2020)
@@
However, in principle, such a post-processing search for a good CV should not be necessary since the committor has been argued to be the best possible one-dimensional reaction coordinate. Berezhkovskii and Szabo (2005); Peters and Trout (2006); Ma and Dinner (2005); Li and Ma (2014); He, Chipot, and Roux (2022) Unfortunately, directly using $q(\mathbf{x})$ as a CV is unpractical, if not impossible, in an enhanced sampling context. Indeed, in the basins, where $q(\mathbf{x})$ is either $\approx 0$ ($\mathbf{x}\in A$) or $\approx 1$ ($\mathbf{x}\in B$), different configurations are distinguished only by very tiny numerical variations of $q(\mathbf{x})$, thus leading to numerical difficulties. Furthermore, another source of numerical issues comes from the very sharp behavior of $q(\mathbf{x})$ in the transition state region, Khoo, Lu, and Ying (2019) further complicating its use as a CV (see Fig. 1A and SI, Fig. S1).
However, a convenient solution can be found if we recall that in Ref. 1 $q(\mathbf{x})$ was expressed as a neural network whose output $z(\mathbf{x})$ was fed to a step-like activation function $\sigma$ to facilitate imposing the right functional form (see Fig. 1A)
@@
The output of the network $z(\mathbf{x})$ encodes the same information as $q(\mathbf{x})$ but is smoothly varying and can be used in practice as an efficient CV. This observation is crucial to the success of the improved approach we present here, which makes the method of Ref. 1 an effective and semi-automatic way to perform free energy calculations. Indeed, we do not wait for the end of the iterative process to perform an additional free energy calculation, but we add to $V_{K}$ an On-the-fly Probability Enhanced Sampling Invernizzi and Parrinello (2020) (OPES) bias that uses $z(\mathbf{x})$ as a CV in the iterative procedure itself and it’s aimed at filling the energy landscape to promote transitions (see Fig. 1B). We shall refer to this combined bias as OPES+$V_{K}$, and we show that this, alongside a few technical improvements, accelerates the convergence of the iterative procedure and leads to a balanced sampling, in which configurations belonging to the metastable states and to the transition state ensemble are equally well sampled, and no additional calculations are needed to obtain a converged free energy estimate (see Fig. 1B).
For illustrative purposes, we first demonstrate the workings of our approach on some simple but informative systems like some two-dimensional models and the classical example of alanine dipeptide in a vacuum. We then move to more complex systems, studying the folding of the chignolin protein in water and the interaction between an organic ligand and a calyxarene, which is a simplified but non-trivial model for drug-protein interaction. In this last example, different binding pathways are possible, and the ligand visits several intermediate states on its way to the binding pocket. This demonstrates the generality of our approach, which does not require assuming that there is only one reaction pathway. A full description of the methods, a summary of the procedure, and other technical details are provided in Sec. III and in the SI.
@@ II.1 Müller potential
Performing OPES+$V_{K}$ simulations allows covering the whole relevant phase space from the metastable states through the TS, as shown in Fig. 2A. Moreover, to achieve this better sampling, it is not necessary to wait until a converged model is available, as good sampling is obtained even from the very first iteration (see SI, Fig. S2) in which the model is not much more than a classifier trained to satisfy the boundary conditions $q({\mathbf{x}}_{A})=0$ and $q({\mathbf{x}}_{B})=1$ (see Sec. III.2). However, this result should come as little surprise, as machine-learning classifier-like CVs have already proved effective in driving enhanced sampling simulations with OPES. Bonati, Rizzi, and Parrinello (2020); Trizio and Parrinello (2021); Ray, Trizio, and Parrinello (2023) In contrast, using only $V_{K}$ in Ref. 1, the sampling was unbalanced toward the TS region and often heavily dependent on the basin from which the simulation was started. Another advantage in terms of sampling comes from the action of the Kolmogorov bias $V_{K}$, which favors extensive sampling of the TS region. This is evident if one compares the distribution along $z$ of the points sampled with a standard OPES simulation and one performed by combining OPES and $V_{K}$ (see Fig. 2B). From the two distributions, it can be clearly appreciated the different emphasis that the two simulations put on the TS region. When only OPES is used, most of the time is spent in the metastable basins. On the other hand, with the combined bias, a large number of TSE configurations is harvested. Such configurations are precious not only for better investigating the TS properties but also for effectively applying the variational principle to optimize the committor.
@@
In addition, a second advantage of this approach is related, besides the quantity of the sampling, to its quality. Thanks to the static nature of $V_{K}$ and the robust OPES framework, Invernizzi and Parrinello (2022) which is designed to quickly converge to a quasi-static regime, very reliable weights can be assigned to the sampled points, thus allowing for a proper reweighting to the true Boltzmann distribution, which, again, has several implications. On the one side, this allows the recovery of the free energy surface (FES), as we show in Fig. 2C. There, we report the free energy profiles along $z$ obtained by reweighting the same two trajectories considered above, compared with an analytical reference. The three curves are almost indistinguishable, thus proving the quality of the reweighting procedure and the (non-obvious) absence of detrimental effects of the $V_{K}$ bias on such free energy calculations.
On the other side, having such reliable weights also for the points in the metastable regions facilitates the convergence of the iterative procedure, at variance with Ref. 1. Indeed, even for this simple example, this approach allows converging the estimate of the committor to the numerical reference almost perfectly in just a couple of iterations (see inset of Fig. 2A).
@@ II.2 Alanine Dipeptide
In most publications devoted to rare event sampling, it is traditional to illustrate the capability of a new method on the conformational equilibrium of alanine dipeptide, and here, we shall stick to this tradition. Moreover, having studied alanine also in our previous work, Kang, Trizio, and Parrinello (2024) we are in the position to asses the improvement in convergence speed due to our improved protocol. There, a good guess for the committor function was obtained from the $6^{\text{th}}$ iteration of the procedure, whereas here, it was already available with half of the iterations (see SI, Table S2). This allowed us not only to efficiently sample the whole relevant phase space (see Fig. 3A) but also to properly identify the TSE (Fig. 3B), which is known to exhibit a linear relation between $\varphi$ and $\theta$.  Bolhuis, Dellago, and Chandler (2000) It is also worth noting that one can still estimate the Kolmogorov distribution $p_{K}$ by reweighting the OPES+$V_{K}(\mathbf{x})$ sampled configurations only with the OPES part of the bias. As discussed in Ref.1, this allows defining in a precise way the TSE.
Also in this case, it is instructive to compare the distribution of points sampled using different bias potentials. In Fig. 3C, we compare the outcome of three different OPES simulations, in one, we use the traditional $\varphi$ and $\psi$ as CVs, in the second the $z(\mathbf{x})$ is used as CV, and finally in the third the action of the $z(\mathbf{x})$ OPES bias is boosted by the Kolmogorov bias $V_{K}(\mathbf{x})$. As seen for the Müller potential, combining the $V_{K}$ potential with OPES systematically enhances the sampling of the TSE while still guaranteeing a thorough sampling of the metastable states. From such data, we can thus recover a correct estimate of the free energy of the system, which is indistinguishable not only from that obtained in the normal OPES run on $z$ but also from the reference simulation using $\varphi$ and $\psi$ (see Fig. 3D), thus increasing our confidence in the method.
@@ II.3 Double Path Potential
Another situation that can occur in practice and which is not straightforward to deal with using other methodologies is the one in which the reaction can follow different pathways Vlugt and Smit (2001); Borrero and Dellago (2016); Bolhuis and Csányi (2018); Mandelli, Hirshberg, and Parrinello (2020); Capelli, Carloni, and Parrinello (2019). A simplified and yet instructive model is the one depicted in Fig. 4, where the system can go from state $A$ to state $B$ following two different pathways, with the upper path characterized by a larger free energy barrier.
With our approach, both paths are sampled starting from the first iteration, thus allowing for a quick convergence of the procedure (see SI, Fig. S4). It is then instructive to focus on the final results reported in Fig. 4. In panel A, we report the converged committor function for the system, showing not only that both paths can be handled at the same time but also that the committor varies more abruptly in the case of the higher barrier, whereas it is more spread for the favored channel. This behavior provides a non-trivial example of why our choice to use the Kolmogorov distribution Kang, Trizio, and Parrinello (2024) $p_{K}$ (see Eq. 7) to identify TSE is to be preferred over the conventional $q\simeq 0.5$ definition. If one considers the $p_{K}$ for this system (see Fig. 4), one can readily see that most of the contribution to the transition flux comes from the lower channel, and only a minor portion goes through the other. In contrast, the $q\simeq 0.5$ criterion would have given the same importance to the two paths and also to much higher energy regions that were actually never sampled, not even in biased simulations (see SI, Fig. S4). The predominance of the lower path is also reflected by the distribution of the sampled points, which are much denser there (see SI, Fig. S4). Nevertheless, after reweighting, an almost perfect 2D FES for the system, including precise estimates of the barriers, can be recovered effortlessly, as shown in Fig. 4C.
@@ II.4 Chignolin: Protein Folding
A more challenging test for our method is the study of the folding of the solvated chignolin protein into a stable hairpin structure. Conveniently, for this system, results from long unbiased MD simulations are also available, Lindorff-Larsen et al. (2011) providing a reference for the folding energy and the overall dynamics of the process. In our previous work, Kang, Trizio, and Parrinello (2024) we already studied this system, characterizing the TSE committor and identifying two possible transition routes. There, we used as descriptors the 45 distances between the alpha carbon atoms, which were able to describe well the protein backbone. Here, we make use of the greater efficiency of our approach to enlarge the number of descriptors, using the 210 distances selected in Ref. 19 that also include the interaction of the backbone with the side chains. As discussed in Ref. 1, better variational flexibility typically leads to more accurate results and to a lower value of the variational bound $K_{m}$. Here, in fact, on the same dataset, the minimum goes from the $K_{m}\approx 19$ to the much lower bound of $K_{m}\approx 2.2$ (see also SI, Sec S4).
@@
In Ref. 1, we have shown that the properties of the transition state ensemble could be analyzed in great and illuminating detail with a combination of k-medoid clustering Schubert and Lenssen (2022) and feature relevance analysis. Pizarroso, Portela, and Muñoz (2022) Here, as we get a thorough sampling of the entire reaction path, we extend this analysis to understand the full folding mechanism.
To do so, we take advantage of the fact that $q(\mathbf{x})$, or equivalently, for simplicity, the smoother $z(\mathbf{x})$, provide a natural ordering in the configurations sampled, that can be used to order them as one goes from $q(\mathbf{x}\in A)=0$ to $q(\mathbf{x}\in B)=1$. Thus, we sliced the whole trajectory into 10 bins along the whole $z$ range (i.e., $-6.5<z<6.5$) and into 10 more focused on the TS region (i.e., $-1.5<z<1.5$), and for each, we identified the most representative configurations through the k-medoid analysis. Schubert and Lenssen (2022) Then, we performed a LASSO-based Novelli et al. (2022) feature relevance analysis to identify which descriptors were more relevant for each bin, revealing that the orientation of the side chains does play a role in the formation of the hydrogen bond network that holds the protein together and regulate steric effects. In Fig. 5, we show the superimposition of several configurations belonging to the most representative bins across the (un)folding process.
A first general observation is that, as the folding proceeds, the entropy within the bins decreases, and the corresponding configurations are more similar to each other. Interestingly, an explanation can be found guided by the LASSO results (see SI, Table S5), which focus the attention on the sidechains of residues Tyr2 and Trp9 in the folded configurations (see Fig. 5 1 and 2). Indeed, moving from frame 1 to 2, the stable folded structure starts to be dismantled due to the misalignment of the side groups of Tyr2 and Trp9, which destroys the hydrophobic $\pi$-$\pi$ interaction between their aromatic rings and owing to the rotation of the side chain of Tyr1, which induces the breaking of the H-bond that locks the end of the main chain.

### v4

# Everything everywhere all at once: a probability-based enhanced sampling approach to rare events

@@ Abstract
The problem of studying rare events is central to many areas of computer simulations. We recently proposed an approach to solving this problem that passes through the computation of the committor function, showing how it can be iteratively computed in a variational way while efficiently sampling the transition state ensemble. Here, we greatly ameliorate this procedure by combining it with a metadynamics-like enhanced sampling approach in which a logarithmic function of the committor is used as a collective variable. This procedure leads to an accurate sampling of the free energy surface in which transition states and metastable basins are studied with the same thoroughness. We show that our approach can be used in cases with the possibility of competing reactive paths and metastable intermediates. In addition, we demonstrate how physical insights can be obtained from the optimized committor model and the sampled data, thus providing a full characterization of the rare event under study.
@@ I Introduction
In the last few decades, our group has contributed to this collective effort and very recently has proposed a new practical method to address the rare event problem, which is based on computing the committor function $q(\mathbf{x})$ and using it as a sampling tool, which we summarize in Section IV.1Kang, Trizio, and Parrinello (2024) We recall here that the committor is a function of the atomic coordinates $\mathbf{x}$ which, given two metastable states $A$ and $B$, is defined as the probability that a trajectory started at $\mathbf{x}$ ends in $B$ without having first passed by $A$Weinan and Vanden-Eijnden (2010) The committor function is arguably the most precisely defined way of describing rare events and is a quantity that remains well-defined even if the transition from $A$ to $B$ follows different competing pathways. As the committor has so many attractive features, much work has been done to determine it. Bolhuis, Dellago, and Chandler (2000); Ma and Dinner (2005); Berezhkovskii and Szabo (2005); Peters and Trout (2006); E. and Vanden-Eijnden (2006); Rotskoff, Mitchell, and Vanden-Eijnden (2022); Chen et al. (2023); Chen, Roux, and Chipot (2023); Li, Lin, and Ren (2019); Jung et al. (2023); Mitchell and Rotskoff (2024); He, Chipot, and Roux (2022); Khoo, Lu, and Ying (2019)
@@
where the average $\langle\cdot\rangle_{U(\mathbf{x})}$ is over the Boltzmann distribution associated with the interatomic potential $U(\mathbf{x})$ at inverse temperature $\beta$, and $\nabla_{\mathbf{u}}$ denotes the gradient with respect to the mass-weighted coordinates.
In Ref. 3, we computed the variational minimum of $K[q(\mathbf{x})]$ using a self-consistent iterative procedure that to be started, only requires that the initial and final states are known. Key to the success of this approach has been the use of a bias potential $V_{K}$ that is a functional of the committor
@@
which is able to focus sampling on the otherwise elusive transition state region (see Figure 1). This approach has proven to be highly successful in finding and extensively sampling the transition state ensemble (TSE). However, additional simulations were needed to evaluate the system free energy since $V_{K}(\mathbf{x})$ disproportionately favors sampling of the transition state ensemble region. This additional calculation can be performed, for instance, using the data collected during the iterative process to generate an efficient collective variable Bonati, Piccini, and Parrinello (2021); Ray, Trizio, and Parrinello (2023) (CV) and then performing a metadynamics-like calculation using this CV. Laio and Parrinello (2002); Invernizzi and Parrinello (2020); Trizio et al. (2024)
@@
However, in principle, such a post-processing search for a good CV should not be necessary since the committor has been argued to be the best possible one-dimensional reaction coordinate. Berezhkovskii and Szabo (2005); Peters and Trout (2006); Ma and Dinner (2005); Li and Ma (2014); He, Chipot, and Roux (2022) Unfortunately, directly using $q(\mathbf{x})$ as a CV is unpractical, if not impossible, in an enhanced sampling context. Indeed, in the basins, where $q(\mathbf{x})$ is either $\approx 0$ ($\mathbf{x}\in A$) or $\approx 1$ ($\mathbf{x}\in B$), different configurations are distinguished only by very tiny numerical variations of $q(\mathbf{x})$, thus leading to numerical difficulties. Furthermore, another source of numerical issues comes from the very sharp behavior of $q(\mathbf{x})$ in the transition state region, Khoo, Lu, and Ying (2019) further complicating its use as a CV (see Figure 1A and Supplementary Figure S1).
However, a convenient solution can be found if we recall that in Ref. 3 $q(\mathbf{x})$ was expressed as a neural network whose output $z(\mathbf{x})$ was fed to a step-like activation function $\sigma$ to facilitate imposing the right functional form (see Figure 1A)
@@
The output of the network $z(\mathbf{x})$ encodes the same information as $q(\mathbf{x})$ but is smoothly varying and can be used in practice as an efficient CV. This observation is crucial to the success of the improved approach we present here, which makes the method of Ref. 3 an effective and semi-automatic way to perform free energy calculations. Indeed, we do not wait for the end of the iterative process to perform an additional free energy calculation, but, during the whole iterative procedure, we simultaneously apply the $V_{K}$ potential and an On-the-fly Probability Enhanced Sampling Invernizzi and Parrinello (2020); Trizio et al. (2024) (OPES) bias that uses $z(\mathbf{x})$ as a CV that it’s aimed at filling the energy landscape to promote transitions (see Figure 1B). We shall refer to this combined bias as OPES+$V_{K}$, and we show that this, alongside a few technical improvements, accelerates the convergence of the iterative procedure and leads to a balanced sampling, in which configurations belonging to the metastable states and to the transition state ensemble are equally well sampled, and no additional calculations are needed to obtain a converged free energy estimate upon reweighting of the sampled configurations by the total bias (see Figure 1B).
For illustrative purposes, we first demonstrate the workings of our approach on some simple but informative systems like some two-dimensional models and the classical example of alanine dipeptide in a vacuum. We then move to more complex systems, studying the folding of the chignolin protein in water and the interaction between an organic ligand and a calyxarene, which is a simplified but non-trivial model for drug-protein interaction. In this last example, different binding pathways are possible, and the ligand visits several intermediate states on its way to the binding pocket. This demonstrates the generality of our approach, which does not require assuming that there is only one reaction pathway. A full description of the methods, a summary of the procedure, and other technical details are provided in Section IV and in the Supporting Information.
@@ II.1 Müller potential
Performing OPES+$V_{K}$ simulations allows covering the whole relevant phase space from the metastable states through the TS, as shown in Figure 2A. Moreover, to achieve this better sampling, it is not necessary to wait until a converged model is available, as good sampling is obtained even from the very first iteration (see Supplementary Figure S2) in which the model is not much more than a classifier trained to satisfy the boundary conditions $q({\mathbf{x}}_{A})=0$ and $q({\mathbf{x}}_{B})=1$ (see Section IV.2). However, this result should not surprise, as machine-learning classifier-like CVs have already proved effective in driving enhanced sampling simulations with OPES. Bonati, Rizzi, and Parrinello (2020); Trizio and Parrinello (2021); Ray, Trizio, and Parrinello (2023) In contrast, using only $V_{K}$ in Ref. 3, the sampling was unbalanced toward the TS region and often heavily dependent on the basin from which the simulation was started. Another advantage in terms of sampling comes from the action of the Kolmogorov bias $V_{K}$, which favors extensive sampling of the TS region. This is evident if one compares the distribution along $z$ of the points sampled with a standard OPES simulation and one performed by combining OPES and $V_{K}$ (see Figure 2B). From the two distributions, it can be clearly appreciated the different emphasis that the two simulations put on the TS region. When only OPES is used, most of the time is spent in the metastable basins. On the other hand, with the combined bias, a large number of TSE configurations is harvested. Such configurations are precious not only for better investigating the TS properties but also for effectively applying the variational principle to optimize the committor.
@@
In addition, a second advantage of this approach is related, besides the quantity of the sampling, to its quality. Thanks to the static nature of $V_{K}$ and the robust OPES framework, Invernizzi and Parrinello (2022); Trizio et al. (2024) which is designed to quickly converge to a quasi-static regime, very reliable weights can be assigned to the sampled points, thus allowing for a proper reweighting to the true Boltzmann distribution, which, again, has several implications. On the one side, this allows the recovery of the free energy surface (FES), as we show in Figure 2C. There, we report the free energy profiles along $z$ obtained by reweighting the same two trajectories considered above, compared with an analytical reference. The three curves are almost indistinguishable, thus proving the quality of the reweighting procedure and the (non-obvious) absence of detrimental effects of the $V_{K}$ bias on such free energy calculations.
On the other side, having such reliable weights also for the points in the metastable regions facilitates the convergence of the iterative procedure, at variance with Ref. 3. Indeed, even for this simple example, this approach allows converging the estimate of the committor to the numerical reference almost perfectly in just two iterations (see inset of Figure 2A).
@@ II.2 Alanine Dipeptide
In most publications devoted to rare event sampling, it is traditional to illustrate the capability of a new method on the conformational equilibrium of alanine dipeptide, and here, we shall stick to this tradition. Moreover, having studied alanine also in our previous work, Kang, Trizio, and Parrinello (2024) we are in the position to assess the improvement in convergence speed due to our improved protocol. There, a good guess for the committor function was obtained from the $6^{\text{th}}$ iteration of the procedure, whereas here, it was already available with half of the iterations (see Supplementary Table S2). This allowed us not only to efficiently sample the whole relevant phase space (see Figure 3A) but also to properly identify the TSE (Figure 3B), which is known to exhibit a linear relation between $\varphi$ and $\theta$.  Bolhuis, Dellago, and Chandler (2000) It is also worth noting that one can still estimate the Kolmogorov distribution $p_{K}$ by reweighting the OPES+$V_{K}(\mathbf{x})$ sampled configurations only with the OPES part of the bias. As discussed in Ref.3, this allows defining in a precise way the TSE.
Also in this case, it is instructive to compare the distribution of points sampled using different bias potentials. In Figure 3C, we compare the outcome of three different OPES simulations, in one, we use the traditional $\varphi$ and $\psi$ as CVs, in the second the $z(\mathbf{x})$ is used as CV, and finally in the third the action of the $z(\mathbf{x})$ OPES bias is boosted by the Kolmogorov bias $V_{K}(\mathbf{x})$. As seen for the Müller potential, combining the $V_{K}$ potential with OPES systematically enhances the sampling of the TSE while still guaranteeing a thorough sampling of the metastable states. From such data, we can thus correctly estimate the free energy of the system, which is indistinguishable not only from that obtained in the normal OPES run on $z$ but also from the reference simulation using $\varphi$ and $\psi$ (see Figure 3D), thus increasing our confidence in the method.
@@ II.3 Double Path Potential
Another situation that can occur in practice and which is not straightforward to deal with using other methodologies is the one in which the reaction can follow different pathways Vlugt and Smit (2001); Borrero and Dellago (2016); Bolhuis and Csányi (2018); Mandelli, Hirshberg, and Parrinello (2020); Capelli, Carloni, and Parrinello (2019). A simplified and yet instructive model is the one depicted in Figure 4, where the system can go from state $A$ to state $B$ following two different pathways, with the upper path characterized by a larger free energy barrier.
With our approach, both paths are sampled starting from the first iteration, thus allowing for a quick convergence of the procedure (see Supplementary Figure S4). It is then instructive to focus on the final results reported in Figure 4. In panel A, we report the converged committor function for the system, showing not only that both paths can be handled at the same time but also that the committor varies more abruptly in the case of the higher barrier, whereas it is more spread for the favored channel. This behavior provides a non-trivial example of why our choice to use the Kolmogorov distribution Kang, Trizio, and Parrinello (2024) $p_{K}$ (see Eq. 7) to identify TSE is to be preferred over the conventional $q\simeq 0.5$ definition. If one considers the $p_{K}$ for this system (see Figure 4), one can readily see that most of the contribution to the transition flux comes from the lower channel, and only a minor portion goes through the other. In contrast, the $q\simeq 0.5$ criterion would have given the same importance to the two paths and also to much higher energy regions that were actually never sampled, not even in biased simulations (see Supplementary Figure S4). The predominance of the lower path is also reflected by the distribution of the sampled points, which are much denser there (see Supplementary Figure S4). Eventually, in more complex cases, a more homogeneous sampling could also be achieved by biasing additional orthogonal CVs that could distinguish them, which is also straightforward in the OPES framework (see SectionIV.3). Nevertheless, in this case, after reweighting, an almost perfect 2D FES for the system, including precise estimates of the barriers, can be recovered effortlessly, as shown in Figure 4C.
@@ II.4 Chignolin: Protein Folding
A more challenging test for our method is the study of the folding of the solvated chignolin protein into a stable hairpin structure. Conveniently, for this system, results from long unbiased MD simulations are also available, Lindorff-Larsen et al. (2011) providing a reference for the folding energy and the overall dynamics of the process. In our previous work, Kang, Trizio, and Parrinello (2024) we already studied this system, characterizing the TSE committor and identifying two possible transition routes. There, we used as descriptors the 45 distances between the alpha carbon atoms, which were able to describe well the protein backbone. Here, we make use of the greater efficiency of our approach to enlarge the number of descriptors, using the 210 distances selected in Ref. 19 that also include the interaction of the backbone with the side chains. As discussed in Ref. 3, better variational flexibility typically leads to more accurate results and to a lower value of the variational bound $K_{m}$. Here, in fact, on the same dataset, the minimum goes from the $K_{m}\approx 19$ to the much lower bound of $K_{m}\approx 2.2$ (see also Supplementary Section S4).
@@
In Ref. 3, we have shown that the properties of the transition state ensemble could be analyzed in great and illuminating detail with a combination of k-medoid clustering Schubert and Lenssen (2022) and feature relevance analysis. Pizarroso, Portela, and Muñoz (2022) Here, as we get a thorough sampling of the entire reaction path, we extend this analysis to understand the full folding mechanism.
To do so, we take advantage of the fact that $q(\mathbf{x})$, or equivalently, for simplicity, the smoother $z(\mathbf{x})$, provide a natural ordering in the configurations sampled, that can be used to order them as one goes from $q(\mathbf{x}\in A)=0$ to $q(\mathbf{x}\in B)=1$. Thus, we sliced the whole trajectory into 10 bins along the whole $z$ range (that is, $-6.5<z<6.5$) and into 10 more focused on the TS region (that is, $-1.5<z<1.5$), and for each, we identified the most representative configurations through the k-medoid analysis. Schubert and Lenssen (2022) Then, we performed a LASSO-based Novelli et al. (2022) feature relevance analysis to identify which descriptors were more relevant for each bin, revealing that the orientation of the side chains does play a role in the formation of the hydrogen bond network that holds the protein together and regulate steric effects. In Figure 5, we show the superimposition of several configurations belonging to the most representative bins across the (un)folding process.
A first general observation is that, as the folding proceeds, the entropy within the bins decreases, and the corresponding configurations are more similar to each other. Interestingly, an explanation can be found guided by the LASSO results (see Supplementary Table S5), which focus the attention on the sidechains of residues Tyr2 and Trp9 in the folded configurations (see Figure 5 1 and 2). Indeed, moving from frame 1 to 2, the stable folded structure starts to be dismantled due to the misalignment of the side groups of Tyr2 and Trp9, which destroys the hydrophobic $\pi$-$\pi$ interaction between their aromatic rings and owing to the rotation of the side chain of Tyr1, which induces the breaking of the H-bond that locks the end of the main chain.